# Novel approach for ranking DEMs: Copernicus DEM improves one arc second open global topography

### 01 Dec 2023 v5

@@ Abstract
Cite as: Bielski, C.; López-Vázquez, C.; Grohmann, C.H.; Guth. P.L. and the TMSG DEMIX Working Group, 2023. Novel approach for ranking DEMs: Copernicus DEM improves one arc second open global topography. Submitted to IEEE Transactions on Geoscience & Remote Sensing.
@@
url]https://ceos.org/ourwork/workinggroups/wgcv/current-activites/
@@ 2.1 Test areas and DEMIX Tiling
For each of the test areas, one or more DEMIX tiles were extracted. A DEMIX tile covers approximately 10 km x 10 km in size defined on a geographic latitude/longitude grid. This grid covers the entire globe and was developed in the context of DEMIX to allow anyone to use a consistent sampling. The DEMIX tiles provide a global tessellation at a scale fine enough to produce locally significant results, coarse enough to keep the total number of tiles manageable, and nearly equal in size to allow for statistical aggregation and comparative analysis. Technical details, naming convention and grid definition files can be found in Guth et al. (2023b).

### 21 Feb 2024 v6

@@ Abstract
Cite as: Bielski, C.; López-Vázquez, C.; Grohmann, C.H.; Guth. P.L.; Hawker, L.; Gesch, D.; Trevisani, S.; Herrera-Cruz, V.; Riazanoff, S.; Corseaux, A.; Reuter, H.; Strobl, P., 2024. Novel approach for ranking DEMs: Copernicus DEM improves one arc second open global topography. IEEE Transactions on Geoscience & Remote Sensing. https://doi.org/10.1109/TGRS.2024.3368015
@@
url]https://orcid.org/$0$-0002-8317-7084
url]https://orcid.org/$0$-0002-8992-4933
url]https://orcid.org/$0$-0001-8436-7798
url]https://orcid.org/$0$-0003-1746-1471
url]https://orcid.org/$0$-0001-6399-7299
url]https://orcid.org/0009-0009-4960-7472
url]https://orcid.org/$0$-0001-6336-7801
url]https://orcid.org/$0$-0003-2733-1822
@@ 2.1 Test areas and DEMIX Tiling
For each of the test areas, one or more DEMIX tiles were extracted. A DEMIX tile covers approximately 10 km x 10 km in size defined on a geographic latitude/longitude grid. This grid covers the entire globe and was developed in the context of DEMIX to allow anyone to use a consistent sampling. The DEMIX tiles provide a global tessellation at a scale fine enough to produce locally significant results, coarse enough to keep the total number of tiles manageable, and nearly equal in size to allow for statistical aggregation and comparative analysis. Technical details, naming convention and grid definition files can be found in Guth et al. (2023c).
@@ 4.2.2 Contest Criteria Ranking Requirements
4.3 Practical Application – Hydrology
Most users want to use DEMs for a wide range of practical applications, ranging from modeling landslide susceptibility, making flood models, to creating hillshade base maps. Statistical measures such as the 15 criteria used in our rankings do not always intuitively relate to what users want to apply, and most comparisons in the scientific literature use only a few specific metrics comparing elevations to reference data. We have argued that our combined assessment of elevation, slope, and roughness of the DEMs results in a more comprehensive and broadly applicable ranking of global DEMs compared to using only elevation metrics focusing on small regions.
[…]
Table 5 shows the four categories resulting from the analysis, and how the hit percentage (Correct channel / (Correct channel + Underprediction)) can be used to assess the results Stephens et al. (2013). In figures 10 and 11 we use a miss percentage (100 -- hit percentage) where lower is better which matches the criteria used in the wine contest and Figures 6 and 8.
Table 5: Possible outcomes of the binary pattern measures computed between the pixels in the reference channel dataset and the six global DEMs.
Test DEM channel Test DEM inter-channel
Reference DEM Correct channel Underprediction
channel (misses)
Reference DEM Overpredictions Correct inter-channel
inter-channel (false positive)
Figure 10 shows the results of comparing the river networks produced using the six global DEMs to the reference DTM and plotting the channel pixel miss percentage (low score is better). The relative order to the outcomes and the separations between the six global DEMs are constant for all 35 tiles. Differences among the tiles reflect the different terrain characteristics.
Figure 10: Channel location comparison computed as the channel miss percentage for 35 DEMIX tiles in the Madrid (Spain) test area. A low score is better: 0% means the channels derived from the candidate DEM perfectly match the reference DTM, and 100% means they do not match at all.
A visual representation of the river network based comparison is shown for three tiles in Figure 11. The top plot (CopDEM) shows the consistently lowest miss percentages (best depiction of the channel network), and the bottom plot (ASTER) presents an example of the DEM which consistently has the highest miss percentage (worst channel network). The easternmost tile, N40RW004B, is largely urban, and has the largest miss percentage compared to the reference DTM. Flat areas are a known problem in channel extraction Zhang et al. (2017), and all DEMs are not good as seen in the urban tile on the right and the main river valley in the center tile. Even in the hilly areas, the ASTER DEM tiles that are shown miss more of the small order streams. In addition to the channel networks, the hillshade from ASTER is much rougher and does not show the ridges and valleys as well as that from CopDEM.
Figure 11: Three tiles in the Madrid (Spain) test area, showing the miss percentages (low score best) for the channel networks on hillshades for CopDEM (top) and ASTER (bottom).
The global DEM intercomparison that produced the rankings in this paper did not directly consider hydrology. The example presented in this section based on channel networks in the Madrid area closely matches the rankings of the six global DEMs in our holistic set of elevation, slope and roughness criteria. A qualitative assessment of hillshade maps for several DEMIX tiles had similar results and ranking of the six global DEMs Guth et al. (2023a). Both the channel network and hillshade results are based on significantly smaller sample sizes, but we find the outcomes support the wine contest DEM intercomparison rankings (see figure 6) which we expect to be generally valid for a broad range of applications. This hydrology experiment verifies the applicability of the global wine contest rankings beyond having better RMSE or LE90 statistical values and the ranking suggests that most practical applications will benefit from using the DEMs recommended by our rankings.
@@ Appendix A Global DEMs charateristics
Supplementary Material
Figure 12: Tile classifications for the 24 DEMIX test areas in our sample, for one landcover classification and 3 geomorphometric landform classifications. Areas are arranged from north to south. This is an expanded version of Figure 3 of the text to include legends.
Figure 13: Difference maps for the 6 global 1” DEMs, by row, for the elevation, slope, and roughness by column. The legends show the percentage of in green that are higher, steeper, and rougher than the reference DTM, and in red that are lower, less steep, and smoother. This is an expanded version of Figure 4 of the text to show difference maps for all 6 DEMs.